The diagonals of parallelogram are 2i and 2j. then area of parallelogram is?
 

Grade:12th pass

6 Answers

sreemoyee kar
6 years ago
→        → →
A=1/2 I a+b I
where a vector and b vector are the diagonals of the parallelogram
=1/2* I 2i*2j I
=1/2*2 I i*j I
=1/2*2*1
=1 unit^2
Pavan Kumar
6 years ago
we cannot determine area of the parallelogram,
    take two lines bisecting each other but change the angle between you will get infinite parallelograms,,,,so by using 2i and 2j we can construct infinite parallelograms each have different areas
Ishan karki
6 years ago
Shreemoy and Pavan I guess here we are looking at area and it is a scalar. So we can just do simple multiply. 
A=1/2 (2*2)
=1/2*4
=2 
 
Correct me with explanation if I am wrong. But I found in severa websites ans as 1 but in my book ans is 2 so literally I am confused.
 
Pavan Kumar
6 years ago
sorry for my previous answer i didnt consider then as vectors,,,
yeah  what you said is correct answer is 2 
  that is for sure                   
prateek
5 years ago
guys what if diagonals of parallelogram are 5i-4j+3k and 3i+ 2j-k
a)171^1/2 units
b)72^1/2  units         
c)171  units
d)72  units
 
shamman date
3 years ago
Steps:
1) The area of a parallelogram having Diagonals as D1 and D2 vector, then there area is given by
1/2 |(D1 x D2)|
 
Here, D 1 = 2i
D2 = 2j
 
Therefore,
Area = 1/2 |(2i x 2j) |
=1/2 * 4
= 2 units.